OpenStudy (anonymous):

What is the best way to classify each equation 6x-14=2(3x-7) A.identity B. contradiction C. neither @shinalcantara

OpenStudy (shinalcantara):

6x - 14 = 2(3x - 7) 6x - 14 = 6x - 14 since right and left of the equation are equal then we can say that they are identical . it's A